a) Calculate the pressure exerted by 0.0153 moles of CO2 gas in a container of 1.467 L at 20.0°C.

 

b) What is the volume of a 15.0 g sample of propane (C3H8) gas at Standard Temperature Pressure?


c) Assuming that the volume of the container in question part (b) above remains constant, what would be the pressure of the propane at room temperature (25 °C)? 



1
Expert's answer
2021-01-20T03:17:09-0500

a) P V = n R T .


P × 1.46 = 0.0153 × 0.0821 × 293


P = 0.0153 × 0.0821 × 293 / 1.46


P = 0.252 atm .




b ) Volume of ethane gas

= ( 15 / 44 ) mole.

Volume of gas = ( 15 / 44 ) × 22.4 = 7.636 L .






c ) p is directly proportional to Temperature at Constant Volume and moles .

P1 / T1 = P2 / T2 at Constant Volume and moles .


1 / 273 = p2 / 298 .


P2 = 298 / 273 = 1.091 atm .